supplied at that price. You will not be graded on any changes you make to this graph. Graph Input Tool 80T Market for Hats Price (Dollars per hat) Supp 24 uantit emanded 500C 56 (Hats) O 48 40 32 e 24 16 Demand 0 50 100 150 200 250 300 350 400 450 500 QUANTITY (Hats)
0 50 100 150 200 250 300 350 400 450 500 QUANTITY (Hats) The equiltbrium price in this market is per hat, and the equlibnum quantitysts boght and sold per month e in this market is S per hat, and the equilibrium quantity is hats bought and sold per month. Complete the following table by indicating at each price whether there is a shortage or surplus in the market, the amount of that shortage or surplus and whether this places upward or downward pressure on prices. Shortage or Surplus Amount Price (Dollars per hat) 48 32 Pressure Shortage or Surplus thesis EC201 Syllabus (1).doc ^ Course Outline docx doc lia-Student Re
Answer #1

1) Equilibrium price = $ 40 because demand and supply curve intersects at this price.

2) 250 hats is the equilibrium quantity

Price Shortage or Surplus Shortage or surplus amount Pressure
48 Surplus Qs - Qd = 300 - 125 = 175 Decrease
32 Shortage Qd - Qs = 375 - 200 = 175 Increase